# Diff-IP2D: Diffusion-Based Hand-Object Interaction Prediction on Egocentric Videos

### v3

@@ IV-B Separate evaluation on hand trajectory and object affordance prediction
We first present the evaluation results on hand trajectory prediction. As Tab. I depicts, our proposed Diff-IP2D outperforms all the baselines on both EK55 and EK100 on WDE and FDE. This is mainly achieved by the devised iter-NAR paradigm of Diff-IP2D alleviating degeneration in AR baselines, as well as the egomotion guidance. The visualization of hand prediction results is shown in Fig. 5. It can be seen that our proposed method can better capture the camera wearer’s intention (such as putting the food in the bowl) and generate more reasonable future trajectories even if lacking past observations for hands (such as reaching out towards the table). Besides, our method can predict a good final hand position although there is a large shift in the early stage (the subfigure in the bottom right corner of Fig. 5), which benefits from our parallel generation with bidirectional constraints. When directly transferring the models trained on Epic-Kitchens to the unseen EG dataset, our method still outperforms the other baselines, improving by 7.0% and 15.3% against the second-best method on WDE and FDE respectively. This reveals the solid generalization capability of our Diff-IP2D across different environments.
The comparison results of object affordance prediction are also shown in Tab. I. Our proposed Diff-IP2D predicts the hotspots with larger SIM, AUC-J, and NSS compared to all the baselines on both Epic-Kitchens data and unseen EG data. Fig. 6 illustrates the predicted contact points with minimum distances to the ground-truth ones. Our proposed method focuses more on objects of interest considering the features of the holistic interaction and potential hand trajectories, and therefore grounds the contact points closer to the ground-truth labels than the counterparts of the baseline.
@@
Figure 6: Visualization of object affordance prediction on Epic-Kitchens. The contact points from ground-truth, Diff-IP2D, and the state-of-the-art baseline OCT [12] are represented by red, white, and blue dots respectively. For a clearer illustration, we additionally put a fixed Gaussian with each contact point as the center.
IV-C Joint evaluation on hand-object interaction prediction
We further compare Diff-IP2D with the other two joint prediction baselines, FHOI [14] and OCT [12], using our proposed object-centric protocol. The video clips containing both ground-truth hand waypoints and contact points are used for evaluation in this experiment. The results are also shown in Tab. I, which indicates that our proposed Diff-IP2D can generate the best object-centric HOI predictions considering the two tasks concurrently on both Epic-Kitchens and unseen EG data. The results also suggest that Diff-IP2D outperforms the baselines on object-centric HOI prediction by focusing more attention on the target objects and predicting reasonable hand trajectories around them.
IV-D Ablation study on egomotion guidance
We provide an ablation study of the egomotion features used to guide MADT denoising on the EK55 and EK100 datasets. Here we replace the MHCA in MADT with a multi-head self-attention module (MHSA) to remove the egomotion guidance while keeping the same parameter number. The experimental results in Tab. II show that the guidance of motion features improves our proposed diffusion-based paradigm noticeably on both hand trajectory prediction and object affordance prediction. This is achieved by narrowing the two gaps caused by 2D-3D ill-posed problem and view difference mentioned in Sec. I. Note that the egomotion guidance is more significant on the EK100 dataset than on the EK55 dataset. The reason could be that EK100 has a larger volume of training data incorporating more diverse egomotion patterns than EK55, leading to a model that can capture human dynamics better.
V Conclusion and insights
In this paper, we propose a novel hand-object interaction prediction method Diff-IP2D. Specifically, we implement the devised denoising diffusion in the latent feature space under our proposed egomotion guidance, and jointly predict future hand trajectories and object affordances with the recovered latents on 2D egocentric videos. According to the experimental results, Diff-IP2D dominates the existing baselines on both off-the-shelf metrics and our new evaluation protocol, suggesting promising applications in artificial intelligence systems. It innovatively learns to recover latent HOI features and forecast future HOI states in parallel with bidirectional constraints, which can serve as a foundation generative paradigm for future works on predictive tasks.

### v4

@@ 1 Motion-Related Gaps and Egomotion Homography
We also provide an illustration comparison between iter-NAR parallel generation and AR generation in Fig. 1. Our proposed iter-NAR paradigm predicts future HOI states in parallel considering bidrectional constraints encompassing both forward and backward constraints within the holistic interaction sequence. It also shifts the limited iterations along the time axis to the sufficient iterations in the diffusion denoising direction (also shown in Fig. IV of the main text). Following the derivation of the previous work DiffuSeq [24] which is used for text generation, here we further mathematically prove that our proposed Diff-IP2D prediction process can be regarded as an iter-NAR process. We first introduce a series of intermediate HOI states $\{\mathbf{F}_{s}^{y}\}_{s=0}^{S}$ decoded from $\{\mathbf{y}_{s}\}_{s=0}^{S}$, where $\mathbf{y}_{s}$ denotes the future part of $\mathbf{z}_{s}$ and $\mathbf{y}_{S}\sim\mathcal{N}(0,\textbf{\text{I}})$. $\mathbf{F}^{x}$ represents the past latent HOI features $F_{\text{seq}}^{\scriptstyle{\text{R}}}$ or $F_{\text{seq}}^{\scriptstyle{\text{L}}}$ from Side-Oriented Fusion Module. $\mathbf{M}$ denotes the egomotion guidance $M_{\text{seq}}$ here and will be extended by other perception information in our future work. Therefore, the inference process of our proposed diffusion-based approach can be formulated as follows: $~{}p_{\scriptscriptstyle\text{Diff-IP2D}}(\mathbf{F}^{y}|\mathbf{F}^{x})$ $=$ ${\sum_{\mathbf{F}^{y}_{S},\ldots,\mathbf{F}^{y}_{1}}\int_{\mathbf{y}_{S},\ldots,\mathbf{y}_{0}}}{p(\mathbf{F}^{y}|\mathbf{y}_{0},\mathbf{F}^{x})}\prod_{s=S,\ldots,1}{p(\mathbf{y}_{s-1}|\mathbf{F}^{y}_{s})p(\mathbf{F}^{y}_{s}|\mathbf{y}_{s},\mathbf{F}^{x},\mathbf{M})}$ $=$ ${\sum_{\mathbf{F}^{y}_{S},\ldots,\mathbf{F}^{y}_{1}}\int_{\mathbf{y}_{S},\ldots,\mathbf{y}_{0}}}p(\mathbf{F}^{y}_{S}|\mathbf{y}_{S},\mathbf{F}^{x})\prod_{s=S-1,\ldots,0}{p(\mathbf{F}^{y}_{s}|\mathbf{y}_{s},\mathbf{F}^{x},\mathbf{M})p(\mathbf{y}_{s}|\mathbf{F}^{y}_{s+1})}$ $=$ $\sum_{\begin{subarray}{c}\mathbf{F}^{y}_{S},\ldots,\mathbf{F}^{y}_{1}\end{subarray}}p(\mathbf{F}^{y}_{S}|\mathbf{y}_{S},\mathbf{F}^{x})\prod_{s=S-1,\ldots,0}{{\int}_{\mathbf{y}_{s}}p(\mathbf{F}^{y}_{s}|\mathbf{y}_{s},\mathbf{F}^{x},\mathbf{M})p(\mathbf{y}_{s}|\mathbf{F}^{y}_{s+1})}.$
[…]
3 Motivation of the Regularization Loss
We propose a regularization term $\mathcal{L}_{\text{reg}}$ in Eq. (4) of the main text for better model optimization. Here we provide more details about the motivation of the regularization loss. In the training process, the tokenizer embeds RGB information and hand-object locations to latent features for the following denoising diffusion process. Here we describe the detailed function of the tokenizer: For each input image, we first exploit a pretrained Temporal Segment Network [32] and extract hand and object RoIAlign [50] features given the detected bounding boxes from [13]. Specifically, the center coordinates of the detected bounding boxes are encoded into the hand and object intermediate features, meaning that the latent features transformed from them encompass the spatial information of hands and objects within each image. After being corrupted to noisy features in the forward process and being denoised in the reverse process, the reconstructed latents are further transformed into locations of future hands and contact points by the predictors, including Hand Trajectory Head and Object Affordance Head. As can be seen, the latents are generated from input HOI states by the tokenizer before the forward process, and are further converted to output HOI states by the predictor after the reverse process. This is why we regarded latents before and after the denoising diffusion process as representing the same “profile” of the input HOI sequence. They both inherently encompass HOI state information in the same interaction duration, and the training process can be further regarded as the predictor distilling HOI state knowledge from the tokenizer. Therefore, we build a closer gradient connection between the tokenizer and the predictor by introducing the regularization term into the training process to enhance the knowledge distillation. Tab. III in the main text presents the improvement in HOI prediction from our proposed regularization strategy.
4 More Details about Datasets and Diff-IP2D Training Configurations
The training sets of EK55 [59] and EK100 [60] contain 8523 and 24148 video clips respectively. Their test sets consist of 1894 and 3513 samples for hand trajectory evaluation, and 241 and 401 samples for object affordance evaluation. In contrast to Epic-Kitchens, the EG dataset [9] offers a smaller data volume, including 1880 training samples, 442 evaluation hand trajectories, and 69 evaluation interaction hotspots. All the training sets are automatically generated following Liu et al. [10]. Note that we exclusively use the test part of the EG dataset to assess generalization ability in the experiments of Sec. IV-B and Sec. IV-C since it contains insufficient training samples for reasonable convergence.
For training Diff-IP2D, we use AdamW optimizer [63] with a learning rate $2$e-$4$. The total loss function is depicted below. The loss weights are initially set as $\lambda_{\text{VLB}}=1$, $\lambda_{\text{traj}}=1$, $\lambda_{\text{aff}}=0.1$, and $\lambda_{\text{reg}}=0.2$. All the networks in Diff-IP2D are trained for $30$ epochs with a batch size of $8$ on 2 A100 GPUs.
$\mathcal{L}_{\text{total}}=\lambda_{\text{VLB}}(\mathcal{L}_{\text{VLB}}^{\scriptstyle{\text{R}}}+\mathcal{L}_{\text{VLB}}^{\scriptstyle{\text{L}}})+\lambda_{\text{traj}}(\mathcal{L}_{\text{traj}}^{\scriptstyle{\text{R}}}+\mathcal{L}_{\text{traj}}^{\scriptstyle{\text{L}}})$
$+\lambda_{\text{aff}}\mathcal{L}_{\text{aff}}+\lambda_{\text{reg}}(\mathcal{L}_{\text{reg}}^{\scriptstyle{\text{R}}}+\mathcal{L}_{\text{reg}}^{\scriptstyle{\text{L}}}).$
5 Additional Experimental Results
5-A Joint Evaluation on the Effect of Egomotion Guidance
We present the supplementary evaluation results in the ablation study on egomotion guidance. Our proposed joint evaluation protocol is applied here to show the positive effect of egomotion guidance for denoising diffusion. As can be seen in Tab. 1, the use of the egomotion features enhances the joint prediction performance of Diff-IP2D on both EK55 and EK100. EK100 has a larger data volume which contains much more human motion patterns than EK55, leading to larger improvement on SIM, AUC-J, and NSS by $3.0\%$, $2.1\%$, and $8.5\%$ respectively.
5-B Ablation Study on Observation Time
We use the EK55 dataset to demonstrate the effect of observation time on HOI prediction performance. We present the change of hand trajectory prediction errors with different input sequence lengths $\{2,4,6,8,10\}$, corresponding to the observation time $\{0.5\,\text{s},1.0\,\text{s},1.5\,\text{s},2.0\,\text{s},2.5\,\text{s}\}$. We first use Diff-IP2D trained with $10$ observation frames to implement zero-shot prediction with different sequence lengths. Fig. 3(a) illustrates that the prediction performance drops significantly when the number of observation frames decreases. In contrast, once our proposed model is trained from scratch with the predefined observation time, it generates plausible prediction results as Fig. 3(b) shows. Especially when the number of observation frames decreases to $4$, our method still outperforms the baseline which is trained from scratch with $10$ observation frames. This demonstrates the strong generation ability of our diffusion-based approach with limited conditions.
(a) Zero-shot prediction
(b) Prediction by models trained from scratch
Figure 3: Ablation study on observation time. The reference line represents the performance of the second-best baseline trained from scratch using $10$ observation frames.
5-C Additional Visualization of Object Affordance Prediction on Epic-Kitchens
We additionally illustrate the predicted contact points with average distances to the ground-truth points on frames of Epic-Kitchens. As Fig. 4 shows, our proposed method still outperforms the second-best baseline considering the center of 10 predicted candidates.
@@ 5-C Additional Visualization of Object Affordance Prediction on Epic-Kitchens
Figure 4: Visualization of object affordance prediction grounded on frames of Epic-Kitchens. The ground-truth contact points are represented by red dots. The contact points predicted by our Diff-IP2D with average distances to the ground-truth points are represented by white dots. The counterparts predicted by OCT [10] are represented by blue dots.
Figure 5: Two additional explanatory cases.
We also provide two cases in which our Diff-IP2D predicts object affordances away from ground truth but more reasonable than the counterparts of the baseline. As Fig. 5 shows, our proposed Diff-IP2D focuses more on “meaningful” parts of objects such as handles even though its prediction has a similar distance away from ground-truth contact points compared to the baseline.
TABLE 2: Comparison of performance on hand trajectory prediction on EgoPAT3D-DT
metric OCT [10] USST [11] Diff-IP2D (ours)
ADE (seen) 0.108 0.082 0.076
FDE (seen) 0.122 0.118 0.112
ADE (unseen) 0.091 0.060 0.055
FDE (unseen) 0.147 0.087 0.083
Final displacement errors of baselines [10, 11] are re-evaluated according to the erratum from Bao et al. [11] in their open-source repository.
5-D Evaluation on EgoPAT3D-DT
We further conduct an additional experiment on a new public dataset EgoPAT3D-DT [61, 11] to compare the performance of our proposed Diff-IP2D and two state-of-the-art baselines, OCT [10] and USST [11]. There is no affordance annotation in EgoPAT3D-DT and thus we only report the results of hand trajectory prediction. Following the previous work [11], we use the fixed ratio 60% to split the past and future sequences at 30 FPS. EgoPAT3D-DT encompasses both seen scenes and unseen scenes, where the unseen scenes are only used for testing. We obtain 6356 training sequences, 846 validation sequences, and 1605 test sequences. As can be seen in Tab. 2, our Diff-IP2D conducted on the iter-NAR paradigm with temporal enhancement outperforms the AR baselines on hand trajectory prediction on the EgoPAT3D-DT dataset. The better performance of our proposed approach on the unseen test scenes also demonstrates its solid generalization ability.
6 Supplementary Technical Details
How the GT future hand trajectories are obtained for training? How good are they? We follow the GT labels of future hand trajectories from Liu et al. [10]. They use a hand-object detector [13] to extract hand bounding boxes for each future image. Each bounding box center is projected to the last observation frame (canvas image) using estimated homography. The homography matrix between the future image and the canvas image is obtained by multiplying sequential homography matrices. The projected hand locations in the canvas image plane constitute future hand waypoints for training and testing our model. The GT annotations are high quality because: (1) the hand-object detector achieves around 90% IOU on egocentric datasets [13], (2) low-quality GT hand trajectories are manually removed by Liu et al. [10], and (3) we have rechecked the quality of GT hand trajectories in this work.
How does the model handle the situation when only one hand is visible in the frames? The above-mentioned hand-object detector identifies the visibility of each hand, providing a side-aware valid mask for our work. During training and testing, we pad zero values to the output features of SOFM at the timestamps when the hand is invisible according to the valid mask. Besides, the mask is also used by MADT while computing self- and cross-attention. Moreover, Hermite spline interpolation is used to fill the missing GT waypoints caused by invisible hands. If one hand, e.g., the left hand, is absent throughout the entire video clip, Diff-IP2D focuses diffusion denoising on the visible side for higher efficiency, as the latent features from our SOFM are side-aware. In these cases, the invisible side cannot be used for both supervision and error calculation.
How are the ground truth 10 contact points obtained? How good are they? In the training process, we use GT contact points from Liu et al. [10], who exploit skin segmentation and fingertip detection to determine fingertip locations within hand-object bounding boxes. Each training video clip has one GT future contact point per valid side, averaged from detected fingertip locations. Diff-IP2D outputs one possible contact point per valid side after each forward process. For testing, we also follow Liu et al.’s [10] evaluation pipeline and use their high-quality GT object hotspot annotations from Amazon Mechanical Turk and rechecked by us. Each video clip has 1-5 GT points as the “contact center” annotated by workers in the canvas frame to generate GT hotspots, manually avoiding the occlusion problem. To generate object affordance predictions, we perform 10 inference samples per valid side and select the predicted contact point closest to the predicted hand trajectories. These 10 selected points represent sampled estimates of possible next-active object locations, which are used to calculate the object hotspot as affordance.